OpenStudy (anonymous):

i am having problems solveing x^2-3x-1=0

OpenStudy (anonymous):

What problem?

OpenStudy (anonymous):

Are you using the Quadratic Formula?

OpenStudy (anonymous):

thats what i dont understand

OpenStudy (anonymous):

did this help?

hero (hero):

meverett, it's the solution. Of course it helps.

OpenStudy (anonymous):

ohhh ... I do not learn by someone giving me a solution ...

hero (hero):

meverett, as I stated earlier, your methods are not that great.

OpenStudy (anonymous):

yes but is that how you simplify it also how did you get it

hero (hero):

You have good intentions meverett, but they are not as effective as you want them to be.

OpenStudy (anonymous):

opps he has a question, maybe time to second think

OpenStudy (anonymous):

Hero is using completing the square method, you can notice that L.H.S of the equation is being reduced to a square form.

OpenStudy (anonymous):

Hero, I think you made a typo in there :) should be x^2 - 3x + 9/4

OpenStudy (anonymous):

and then applying square root gives the solution.

OpenStudy (anonymous):

\[\Large \begin{array}{l} \frac{{ - b \pm \sqrt {{b^2} - 4ac} }}{{2a}}\\ {x^2} - 3x - 1 = 0\\ a = 1,b = - 3,c = - 1\\ \frac{{3 \pm \sqrt {{{( - 3)}^2} - (4 \cdot 1 \cdot - 1)} }}{{2 \cdot 1}}\\ = \frac{{3 \pm \sqrt {9 + 4} }}{2}\\ = \frac{{3 \pm \sqrt {13} }}{2} \end{array}\]

hero (hero):

Yep, you're right jemurray

OpenStudy (anonymous):

Hero, how did we get this?

hero (hero):

x^2-3x-1=0 x^2-3x= 1 x^2-3x + 9/4 = 1 + 9/4 (x-3/2)^2 = 13/4 x - 3/2 = +/-sqrt(13/4) x - 3/2 = +/=sqrt(13)/2 x = +/-sqrt(13)/2 +3/2 x = (+/-sqrt(13) + 3)/2

OpenStudy (anonymous):

4x^2 +x-1=0 will this be 8X^2-11

myininaya (myininaya):

I think we should all show our methods. I don't think its good to tell someone there method is bad. Lets just all share and have fun. ;)

OpenStudy (anonymous):

That's poor clarifying, myininaya.

myininaya (myininaya):

ok so zonk you are asking how to solve 4x^2+x-1=0?

hero (hero):

Sometimes telling someone their method is not that great is a good thing. If you're a teacher, you wouldn't lie to your student and say their method is good when it isn't. Total honesty applies here mya.

myininaya (myininaya):

Well I wasn't using the quadratic formula meverett's method?

myininaya (myininaya):

Thats the only method I see by him and I think it is a good one

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You come across as quite aggressive, condescending and belittling, Hero. If we're being honest. But this is straying further from the topic.

OpenStudy (anonymous):

I agree with mya. The discussion initiated by hero in three forums today has been inappropriate, then he goes on and on and leaves the student hanging, never addressing the question. QED

myininaya (myininaya):

\[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

myininaya (myininaya):

Ok lets all just be cool okay?

hero (hero):

meverett, you were the person whose methods I addressed, so of course you're not too happy with me.

OpenStudy (anonymous):

I agree fsm and we are way off topic I try to motivate not hate ...

OpenStudy (anonymous):

I am not happy not because you disagree, but because being off topic is inappropriate and not helping the student.

myininaya (myininaya):

lets help zonk with his problem

OpenStudy (anonymous):

thanks

myininaya (myininaya):

\[4x^2+x-1=0\] so a=4,b=1,c=-1

myininaya (myininaya):

we plugged this into the formula we have above

myininaya (myininaya):

\[x=\frac{-(1) \pm \sqrt{(1)^2-4(4)(-1)}}{2(4)}\]

myininaya (myininaya):

We could show this by completing the square.

myininaya (myininaya):

How about you show him fool? The floor is yours.

myininaya (myininaya):

\[x=\frac{-1 \pm \sqrt{1-16(-1)}}{8}=\frac{-1 \pm \sqrt{1+16}}{8}=\frac{-1 \pm \sqrt{17}}{8}\]

myininaya (myininaya):

And I think fool wants to show you how to derive the quad formula

OpenStudy (anonymous):

was that the end

hero (hero):

lol

myininaya (myininaya):

zonk would you like me to show you how the quad formula is derived?

hero (hero):

How did this turn into a discussion about how to derive the quad formula. Seems like mya and fool are going at it head to head to me

OpenStudy (anonymous):

Me encanta el chisme.

OpenStudy (anonymous):

Let me derive in steps: We want to solve the equation \(ax^2+bx+c=0\) where \( a \ne 0 \).

OpenStudy (anonymous):

The usual argument starts by *dividing* by \(a \) let multiply both sides by \( 4a \)

OpenStudy (anonymous):

We will get \( 4a^2x^2 +4abx+4ac=0 \)

myininaya (myininaya):

\[ax^2+bx+c=0\] a does not equal to zero Divide both sides by a \[x^2+\frac{b}{a}x+\frac{c}{a}=0\] Subtract c/a on both sides \[x^2+\frac{b}{a}x=-\frac{c}{a}\] Now we add (b/2a)^2 on both sides (this completes the square on the left hand side \[x^2+\frac{b}{a}x+(\frac{b}{2a})^2=\frac{-c}{a}+(\frac{b}{2a})^2\] \[(x+\frac{b}{2a})^2=\frac{-c}{a}+\frac{b^2}{4a^2}\] Now the goal is to solve for x So we need to take square root of both sides \[x+\frac{b}{2a}=\pm \sqrt{\frac{-c}{a}+\frac{b^2}{4a^2}}\] Now subtract b/2a on both sides \[x=-\frac{b}{2a} \pm \sqrt{\frac{-c}{a}+\frac{b^2}{4a^2}}\] Now let' s combine the fractions inside the square root \[x=\frac{-b}{2a} \pm \sqrt{\frac{-4ac+b^2}{4a^2}}\] Now we can write \[x=\frac{-b}{2a} \pm \frac{\sqrt{b^2-4ac}}{2a}\] We can combine the fractions \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

OMG

OpenStudy (anonymous):

okay now, \( 4a^2x^2+4abx \) is almost the square of \( 2ax+b \). Precisely, \( 4a^2x^2+4abx=(2ax+b)^2-b^2 \)

OpenStudy (anonymous):

So, our equation reduces to \( (2ax+b)^2 -b^2+4ac=0 \Rightarrow (2ax+b)^2=b^2-4ac \)

OpenStudy (anonymous):

it's almost done: \( 2ax+b=\pm\sqrt{b^2-4ac} \Rightarrow x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \)

OpenStudy (anonymous):

Nice.

OpenStudy (anonymous):

Notice there is no fraction till the end, that's the beauty of this approach!

OpenStudy (anonymous):

the final question is did you get it zonk ?

OpenStudy (anonymous):

lol, he/she should by now.